# Haoye Tian

### v4

@@ Abstract
Code review, which aims at ensuring the overall quality and reliability of software, is a cornerstone of software development. Unfortunately, while crucial, Code review is a labor-intensive process that the research community is looking to automate. Existing automated methods rely on single input-output generative models and thus generally struggle to emulate the collaborative nature of code review. This work introduces CodeAgent, a novel multi-agent Large Language Model (LLM) system for code review automation. CodeAgent incorporates a supervisory agent, QA-Checker, to ensure that all the agents’ contributions address the initial review question. We evaluated CodeAgent on critical code review tasks: (1) detect inconsistencies between code changes and commit messages, (2) identify vulnerability introductions, (3) validate code style adherence, and (4) suggest code revisions. The results demonstrate CodeAgent’s effectiveness, contributing to a new state-of-the-art in code review automation. Our data and code are publicly available (https://github.com/Code4Agent/codeagent).
@@ 1 Introduction
Code review Bacchelli and Bird (2013); Bosu and Carver (2013); Davila and Nunes (2021) implements a process wherein software maintainers examine and assess code contributions to ensure quality and adherence to coding standards, and identify potential bugs or improvements. In recent literature, various approaches Tufano et al. (2021, 2022) have been proposed to enhance the performance of code review automation. Unfortunately, major approaches in the field ignore a fundamental aspect: the code review process is inherently interactive and collaborative Bacchelli and Bird (2013). Instead, they primarily focus on rewriting and adapting the submitted code Watson et al. (2022); Thongtanunam et al. (2022); Staron et al. (2020). In this respect, an effective approach should not only address how to review the submitted code for some specific needs (e.g., vulnerability detection). Still, other non-negligible aspects of code review should also be considered, like detecting issues in code formatting or inconsistencies in code revisions. However, processing multiple sub-tasks requires interactions among different staff members in a real code review scenario, which makes it chellenging to design a bot that performs code review automatically.
@@
Agent-based systems are a newly emerging paradigm and a computational framework in which autonomous entities (aka agents) interact with each other Li et al. (2023a); Qian et al. (2023); Hong et al. (2023) to perform a task. Agent-based approaches have been proposed to address a spectrum of software engineering tasks Qian et al. (2023); Zhang et al. (2024), moving beyond the conventional single input-output paradigm due to their remarkable ability to simulate and model complex interactions and behaviors within dynamic environments Xi et al. (2023); Wang et al. (2023). Recently, with multi-agent systems, the strengths of various agents are utilized to simulate human-like decision-making processes Du et al. (2023); Liang et al. (2023); Park et al. (2023), improving performance in many tasks Chen et al. (2023); Li et al. (2023b); Hong et al. (2023). This paradigm fits well with the challenge of code review, which involves multiple reviewers interacting based on their diverse skills and roles to achieve a thorough review of code.
@@
This paper. Drawing from the success of agent-based collaboration, we propose a multi-agent-based framework CodeAgent to simulate the dynamics of a collaborative team engaged in the code review process, incorporating various roles such as code change authors, reviewers, and decision makers. In particular, as a major contribution towards CodeAgent, we address the challenge of prompt drifting Zheng et al. (2024); Yang et al. (2024), commonly found in multi-agent systems and Chain-of-Thought (CoT) reasoning. This issue, characterized by conversations that stray from the main topic, highlights the need for strategies to keep discussions on track and coherent Greyling (2023); Chae et al. (2023). This drift, often triggered by the model-inspired tangents or the randomness of Large Language Models (LLMs), necessitates the integration of a supervisory agent. We employ an agent named QA-Checker (for "Question-Answer Checker") that monitors the flow of conversation, ensuring that questions and responses remain relevant and on track, thus maintaining the dialogue’s intended objective. Such an agent not only refines queries but also realigns answers to match the original intent, employing a systematic approach grounded in a mathematical framework.
@@
To evaluate the performance of CodeAgent, we first assess its effectiveness for typical review objectives such as finding vulnerabilities 4.1 and validating the consistency and alignment of the code format 4.2. We also compare CodeAgent with state-of-the-art generic and code-specific language models Feng et al. (2020); OPENAI (2022). Finally, we assess the performance of CodeAgent compared to the state of the art for code revision suggestions Tufano et al. (2021); Thongtanunam et al. (2022); Tufano et al. (2022). Since each of these related works presents a specific dataset, we also employ them toward a fair comparison. At the same time, we also collect pull requests from GitHub, featuring an extensive array of commits, messages, and comments to evaluate advanced capabilities.The experimental results display that CodeAgent significantly outperforms the state-of-the-art, achieving a 41% increase in hit rate for detecting vulnerabilities. CodeAgent also excels in consistency checking and format alignment, outperforming the target models. Finally, CodeAgent showcases its robustness for code revision by demonstrating superior average edit progress.
@@
  • We build a new dataset with about than 3 500 real-world code changes and commit messages. This dataset, which includes all relevant files and details in a standalone manner, is useful for assessing advanced code review tasks such as vulnerability detection, code style detection, and code revision suggestions.
  • We demonstrate the effectiveness of the QA-Checker. This agent monitors the conversation flow to ensure alignment with the original intent, addressing the common prompt drifting issues in multi-agent systems.
Experimental evaluation highlights the performance of CodeAgent: In terms of vulnerability detection, CodeAgent outperforms GPT-4 and CodeBERT by 3 to 7 percentage points in terms of the number of vulnerabilities detected. In terms of format alignment, CodeAgent outperforms ReAct by about 14% in terms of recall of inconsistency detection. On the code revision task, CodeAgent surpasses the state of the art in the software engineering literature: in terms of the Edit Progress metric Zhou et al. (2023), CodeAgent offers an average performance improvement of about 30%.
@@ 2 CodeAgent
This section details the methodology behind our novel CodeAgent framework. We discuss tasks and definition in Sec 2.1, pipeline in Section 2.2, defined role cards in Section 2.3, and the design of the QA-Checker in Sec 2.4.
@@ 2.1 Tasks
CR Zhou et al. (2023): Code revision; this task attempts to automatically suggest rewrites of the code change to address any issue discovered.
@@ 2.2 Pipeline
We defined six characters and four phases for the framework. The roles of the characters are illustrated in Figure 1. Each phase contains multiple conversations, and each conversation happens between agents. The four phases consist of 1) Basic Info Sync, containing the roles of chief executive officer (CEO), chief technology officer (CTO), and Coder to conduct modality and language analysis; 2) Code Review, leveraging the Coder and Reviewer for actual code review (i.e., target sub-tasks); 3) Code Alignment, supporting the Coder and Reviewer to correct the commit by code revision or suggestion feedback to the author; and 4) Document, finalizing by synthesizing the opinions of the CEO, CPO, Coder, and Reviewer to give out the final comments. In addition to six defined roles, the proposed architecture of CodeAgent consists of phase-level and conversation-level components. The waterfall model breaks the code review process at the phase level into four sequential phases. At the conversation level, each phase is divided into atomic conversations. These atomic conversations involve task-oriented role-playing between two agents, promoting collaborative communication. One agent works as an instructor and the other as an assistant. Communication follows an instruction-following style, where agents interact to accomplish a specific subtask within each conversation, and each conversation is supervised by QA-Checker. QA-Checker is used to align the consistency of questions and answers between the instructor and the assistant in a conversation to avoid digression. QA-Checker will be introduced in Section 2.4.
Figure 2 shows an illustrative example of the CodeAgent pipeline. CodeAgent receives the request to do the code review with the submitted commit, commit message, and original files. In the first phase, CEO, CPO, and Coder will cooperate to recognize the modality of input (for example, document, code) and language (for example, Python, Java and Go). In the second phase, with the help of Coder, Reviewer will write an analysis report on consistency analysis, vulnerability analysis, format analysis and suggestions for code revision. Then, in the third phase, according to analysis reports, Coder will align or revise the code if it finds incorrect snippets with the help of Reviewer. $Coder$ cooperates with CTO and CEO to summarize the document and codes about the whole code review in the final phase.
@@ 3 Experimental Setup
We evaluate the performance of CodeAgent in various qualitative and quantitative experiments in nine programming languages, on four metrics. In this section, we will discuss experimental settings, including datasets, metrics, and baselines. For more information, please see Appendix C.
@@ 3.2 Metrics
  • Hit Rate (Rate) We also use hit rate to evaluate the rate of confirmed vulnerable issues out of the found issues by approaches on task VA.
@@ 3.3 State-of-the-Art Tools and Models
Our study evaluates various tools and models for code revision and modeling. Trans-Review Tufano et al. (2021) employs src2abs for code abstraction, effectively reducing vocabulary size. AutoTransform Thongtanunam et al. (2022) uses Byte-Pair Encoding for efficient vocabulary management in pre-review code revisions. T5-Review Tufano et al. (2022) leverages the T5 architecture, emphasizing improvement in code review through pre-training on code and text data. In handling both natural and programming languages, CodeBERT Feng et al. (2020) adopts a bimodal approach, while GraphCodeBERT Guo et al. (2021) incorporates code structure into its modeling. CodeT5 Wang et al. (2021), based on the T5 framework, is optimized for identifier type awareness, aiding in generation-based tasks. Additionally, we compare these tools with GPT OPENAI (2022) by OpenAI, notable for its human-like text generation capabilities in natural language processing. Finally, we involve COT Wei et al. (2022) and ReAct Yao et al. (2022), of which COT is a method where language models are guided to solve complex problems by generating and following a series of intermediate reasoning steps and ReAct synergistically enhances language models by interleaving reasoning and action generation, improving task performance and interpretability across various decision-making and language tasks.

### v5

@@ Abstract
Code review, which aims at ensuring the overall quality and reliability of software, is a cornerstone of software development. Unfortunately, while crucial, Code review is a labor-intensive process that the research community is looking to automate. Existing automated methods rely on single input-output generative models and thus generally struggle to emulate the collaborative nature of code review. This work introduces CodeAgent, a novel multi-agent Large Language Model (LLM) system for code review automation. CodeAgent incorporates a supervisory agent, QA-Checker, to ensure that all the agents’ contributions address the initial review question. We evaluated CodeAgent on critical code review tasks: (1) detect inconsistencies between code changes and commit messages, (2) identify vulnerability introductions, (3) validate code style adherence, and (4) suggest code revision. The results demonstrate CodeAgent’s effectiveness, contributing to a new state-of-the-art in code review automation. Our data and code are publicly available (https://github.com/Code4Agent/codeagent).
@@ 1 Introduction
Code review Bacchelli and Bird (2013); Bosu and Carver (2013); Davila and Nunes (2021) implements a process wherein software maintainers examine and assess code contributions to ensure quality and adherence to coding standards, and identify potential bugs or improvements. In recent literature, various approaches Tufano et al. (2021, 2022) have been proposed to enhance the performance of code review automation. Unfortunately, major approaches in the field ignore a fundamental aspect: the code review process is inherently interactive and collaborative Bacchelli and Bird (2013). Instead, they primarily focus on rewriting and adapting the submitted code Watson et al. (2022); Thongtanunam et al. (2022); Staron et al. (2020). In this respect, an effective approach should not only address how to review the submitted code for some specific needs (e.g., vulnerability detection Chakraborty et al. (2021); Yang et al. (2024a)). Still, other non-negligible aspects of code review should also be considered, like detecting issues in code formatting or inconsistencies in code revision Oliveira et al. (2023); Tian et al. (2022); Panthaplackel et al. (2021). However, processing multiple sub-tasks requires interactions among employees in different roles in a real code review scenario, which makes it challenging to design a model that performs code review automatically.
@@
Agent-based systems are an emerging paradigm and a computational framework in which autonomous entities (aka agents) interact with each other Li et al. (2023a); Qian et al. (2023); Hong et al. (2023) to perform a task. Agent-based approaches have been proposed to address a spectrum of software engineering tasks Qian et al. (2023); Zhang et al. (2024); Tang et al. (2023); Tian et al. (2023), moving beyond the conventional single input-output paradigm due to their exceptional ability to simulate and model complex interactions and behaviors in dynamic environments Xi et al. (2023); Yang et al. (2024b); Wang et al. (2023). Recently, multi-agent systems have leveraged the strengths of diverse agents to simulate human-like decision-making processes Du et al. (2023); Liang et al. (2023); Park et al. (2023), leading to enhanced performance across various tasks Chen et al. (2023); Li et al. (2023b); Hong et al. (2023). This paradigm is well-suited to the challenge of code review, where multiple reviewers, each with diverse skills and roles, collaborate to achieve a comprehensive review of the code..
@@
This paper. Drawing from the success of agent-based collaboration, we propose a multi-agent-based framework CodeAgent to simulate the dynamics of a collaborative team engaged in the code review process, incorporating diverse roles such as code change authors, reviewers, and decision makers. In particular, A key contribution of CodeAgent is that we address the challenge of prompt drifting Zheng et al. (2024); Yang et al. (2024c), a common issue in multi-agent systems and Chain-of-Thought (CoT) reasoning. This issue, characterized by conversations that stray from the main topic, highlights the need for strategies to maintain focus and coherence Greyling (2023); Chae et al. (2023). This drift, often triggered by the model-inspired tangents or the randomness of Large Language Models (LLMs), necessitates the integration of a supervisory agent. We employ an agent named QA-Checker (for "Question-Answer Checker") that monitors the conversation flow, ensuring that questions and responses stay relevant and aligned with the dialogue’s intended objective. Such an agent not only refines queries but also realigns answers to match the original intent, employing a systematic approach grounded in a mathematical framework.
@@
To evaluate the performance of CodeAgent, we first assess its effectiveness for typical review objectives such as detecting vulnerabilities 4.1 and validating the consistency and alignment of the code format 4.2. We then compare CodeAgent with state-of-the-art generic and code-specific language models like ChatGPT OPENAI (2022) and CodeBERT Feng et al. (2020). Finally, we assess the performance of CodeAgent compared to the state-of-the-art tools for code revision suggestions Tufano et al. (2021); Thongtanunam et al. (2022); Tufano et al. (2022). Since each of these related works presents a specific dataset, we also employ them toward a fair comparison. Additionally, we also collect pull requests from GitHub, featuring an extensive array of commits, messages, and comments to evaluate advanced capabilities.The experimental results reveal that CodeAgent significantly outperforms the state-of-the-art, achieving a 41% increase in hit rate for detecting vulnerabilities. CodeAgent also excels in consistency checking and format alignment, outperforming the target models. Finally, CodeAgent showcases its robustness for code revision by presenting superior average edit progress.
@@
  • We build a new dataset comprising 3 545 real-world code changes and commit messages. This dataset, which includes all relevant files and details in a self-contained format, is valuable for evaluating advanced code review tasks such as vulnerability detection, code style detection, and code revision suggestions.
  • We demonstrate the effectiveness of the QA-Checker. This agent monitors the conversation flow to ensure alignment with the original intent, effectively addressing the common prompt drifting issues in multi-agent systems.
Experimental evaluation highlights the performance of CodeAgent: In vulnerability detection, CodeAgent outperforms GPT-4 and CodeBERT by 3 to 7 percentage points in terms of the number of vulnerabilities detected. For format alignment, CodeAgent outperforms ReAct by approximately 14% in recall for inconsistency detection. On the code revision task, CodeAgent surpasses the state of the art in software engineering literature, achieving an average performance improvement of about 30% in the Edit Progress metric Zhou et al. (2023).
@@ 2 CodeAgent
This section details the methodology behind our CodeAgent framework. We discuss tasks and definition in Sec 2.1, pipeline in Section 2.2, defined role cards in Section 2.3, and the design of the QA-Checker in Sec 2.4.
@@ 2.1 Tasks
CR Zhou et al. (2023): Code revisions; this task attempts to automatically suggest rewrites of the code change to address any issue discovered.
@@ 2.2 Pipeline
We defined six characters and four phases for the framework. The roles of the characters are illustrated in Figure 1. Each phase contains multiple conversations, and each conversation happens between agents. The four phases consist of 1) Basic Info Sync, containing the roles of chief executive officer (CEO), chief technology officer (CTO), and Coder to conduct modality and language analysis; 2) Code Review, asking the Coder and Reviewer for actual code review (i.e., target sub-tasks); 3) Code Alignment, supporting the Coder and Reviewer to correct the commit through code revision and suggestions to the author; and 4) Document, finalizing by synthesizing the opinions of the CEO, CPO (Chief Product Officer), Coder, and Reviewer to provide the final comments. In addition to six defined roles, the proposed architecture of CodeAgent consists of phase-level and conversation-level components. The waterfall model breaks the code review process at the phase level into four sequential phases. At the conversation level, each phase is divided into atomic conversations. These atomic conversations involve task-oriented role-playing between two agents, promoting collaborative communication. One agent works as an instructor and the other as an assistant. Communication follows an instruction-following style, where agents interact to accomplish a specific subtask within each conversation, and each conversation is supervised by QA-Checker. QA-Checker is used to align the consistency of questions and answers between the instructor and the assistant in a conversation to avoid digression. QA-Checker will be introduced in Section 2.4.
Figure 2 shows an illustrative example of the CodeAgent pipeline. CodeAgent receives the request to do the code review with the submitted commit, commit message, and original files. In the first phase, CEO, CTO, and Coder will cooperate to recognize the modality of input (e.g., document, code) and language (e.g., Python, Java and Go). In the second phase, with the help of Coder, Reviewer will write an analysis report on consistency analysis, vulnerability analysis, format analysis and suggestions for code revision. In the third phase, based on analysis reports, Coder will align or revise the code if any incorrect snippets are identified with assistance from Reviewer. $Coder$ cooperates with CPO and CEO to summarize the document and codes about the whole code review in the final phase.
@@ 3 Experimental Setup
We evaluate the performance of CodeAgent through various qualitative and quantitative experiments across nine programming languages, using four distinct metrics. In this section, we will discuss experimental settings, including datasets, metrics, and baselines. For more information, please see Appendix C.
@@ 3.2 Metrics
  • Hit Rate (Rate) We also use hit rate to evaluate the rate of confirmed vulnerability issues out of the found issues by approaches on task VA.
@@ 3.3 State-of-the-Art Tools and Models
Our study evaluates various tools and models for code revision and modeling. Trans-Review Tufano et al. (2021) employs src2abs for code abstraction, effectively reducing vocabulary size. AutoTransform Thongtanunam et al. (2022) uses Byte-Pair Encoding for efficient vocabulary management in pre-review code revision. T5-Review Tufano et al. (2022) leverages the T5 architecture, emphasizing improvement in code review through pre-training on code and text data. In handling both natural and programming languages, CodeBERT Feng et al. (2020) adopts a bimodal approach, while GraphCodeBERT Guo et al. (2021) incorporates code structure into its modeling. CodeT5 Wang et al. (2021), based on the T5 framework, is optimized for identifier type awareness, aiding in generation-based tasks. Additionally, we compare these tools with GPT OPENAI (2022) by OpenAI, notable for its human-like text generation capabilities in natural language processing. Finally, we involve COT Wei et al. (2022) and ReAct Yao et al. (2022), of which COT is a method where language models are guided to solve complex problems by generating and following a series of intermediate reasoning steps and ReAct synergistically enhances language models by interleaving reasoning and action generation, improving task performance and interpretability across various decision-making and language tasks.